OpenStudy (anonymous):

sinx=3/5 Find sin 2x, cos 2x, and tan 2x from the given information in quadrant I

OpenStudy (cwrw238):

you can find the values of sin , cos and tan x from this right angled triangle: 

OpenStudy (cwrw238):

eg cos x = 4/5

OpenStudy (cwrw238):

now to find sin 2x use the identity sin 2x = 2 sin x cos x and cos 2x = 2 cos^2 x - 2

OpenStudy (anonymous):

how do I find sin 3/5 on the unit cirlcle? is that how I'm supposed to get it?

OpenStudy (cwrw238):

* thats cos 2x = 2 cos^2 x - 1

OpenStudy (cwrw238):

the 3-4-5 right angled triangle gives you sin x = opp/hyp = 3/5

OpenStudy (anonymous):

oh ok

OpenStudy (cwrw238):

to get tan 2x you use the identity tan 2x = sin 2x / cos 2x